# Efficient Bayesian analysis of kilonovae and GRB afterglows with fiesta

### v1

@@ 6 Discussion and Conclusion
The fiesta package provides ML surrogates of physical KN and GRB afterglow models to enable efficient likelihood evaluation when sampling a model posterior from photometric transient data. It offers a flexible API for training these surrogates, enabling the use of effectively three different architectures. Since fiesta utilizes the jax framework, training can be hardware-accelerated by execution on a GPU. Additionally, fiesta includes functionalities to sample the lightcurve posterior with flowMC (Wong et al., 2023a), an adaptive MCMC sampler that uses gradient-based samplers and normalizing flow proposals that can also be run on GPUs.
@@
To demonstrate fiesta’s features, we presented surrogates for the afterglowpy (Ryan et al., 2020) and pyblastafterglow (Nedora et al., 2024) GRB afterglow models, as well as for the KN model from possis (Bulla, 2019, 2023). We find that the surrogates perform well against the respective test data sets, as the prediction error is usually bound within 0.3–0.5 mag. When used during inference, this finite prediction error needs to be offset with a systematic uncertainty $\sigma_{\rm sys}$ that can also be sampled in a flexible manner within fiesta’s implementation. We applied these surrogates to two events with joint KN and GRB afterglow emissions, namely AT2017gfo/GRB170817A and GRB211211A. Using our surrogates, we find similar results to previous analyses of these events, even when using the new pyblastafterglow model for the GRB afterglow. However, our posteriors can be evaluated within minutes, whereas previous analyses, e.g., in Pang et al. (2023); Koehn et al. (2025) could take hours to days, mainly due to the more expensive likelihood evaluation when computing the lightcurve directly from afterglowpy.
@@
Despite the advantages regarding speed when relying on ML surrogates in the likelihood evaluation, this approach also comes with some drawbacks. For one, the parameter ranges in which our surrogates interpolate the training data are fixed, and therefore, it is not possible to extend priors without retraining the surrogates. However, these surrogates can easily be fine-tuned on new training datasets for future use in other places in the parameter space. Further, the prediction error from the surrogate compared to the physical base model could introduce biases in the posteriors. We assessed these potential biases by creating afterglowpy injection-recoveries, where the posterior can be obtained with the surrogate and with the base model. We find that the values for the KL divergence match theoretical expectations within the surrogate uncertainty (Bevins et al., 2025). Further, while we find biases in the recovery of certain parameters from our P-P plots in Fig. 5 and 6, we do not find that this bias is necessarily caused by the use of our surrogates. Still, failure of the surrogate in certain parameter regions can not be excluded, and sanity checks are recommended to validate the results. For instance, the physical base model could be run with the best-fit parameters from the posterior to verify that they actually provide a good fit to the data.
@@
The use of ML techniques for accelerated likelihood evaluation and efficient sampling in Bayesian inference is a common practice in various areas of astrophysics. For instance, the analysis of X-ray spectra has recently been accelerated by the use of advanced stochastic samplers written in jax (Dupourqué et al., 2024), or with the help of ML surrogate models (Barret & Dupourqué, 2024; Tutone et al., 2025; Dupourqué & Barret, 2025). For supernovae, Gaussian process surrogate models were developed for inference in Simongini et al. (2024, 2025). Leeney (2025) created a jax-compatible version of sncosmo Barbary et al. (2025). In fiesta, we focused on surrogates for electromagnetic transients of BNS mergers, but in principle, surrogates for other types of transients could also be incorporated. Additionally, as fiesta surrogates utilize the jax framework, this opens the possibility for combination with other jax software, such as jim (Wong et al., 2023b; Wouters et al., 2024) for the parameter estimation on GW signals from BNS. This would enable fast sampling for the joint analysis of GW signals and electromagnetic counterparts and could be linked further to the inference of the EOS for NS matter through the jester (Wouters et al., 2025) package, which also leverages jax to efficiently evaluate the likelihood of an EOS candidate with a given set of NS observations. Additionally, the backwards compatibility of our surrogates to nmma provides the opportunity to perform efficient Bayesian model comparison to either compare transient types (Breschi et al., 2021; Kunert et al., 2024; Hussenot-Desenonges et al., 2024) or to investigate different systematics in the modeling, e.g., regarding the KN ejecta morphology (Heinzel et al., 2021; King et al., 2025; Hussenot-Desenonges et al., 2025), the KN heating rates and opacities (Tanaka et al., 2020; Bulla, 2023; Sarin & Rosswog, 2024; Brethauer et al., 2024), the GRB jet structure (Kunert et al., 2024; Hayes et al., 2020; Lin et al., 2021) or environment (Aksulu et al., 2022), or the mapping from BNS parameters to ejecta parameters (Ristić et al., 2025). Given the launch of new transient observatories such as the Vera C. Rubin Observatory (Andreoni et al., 2024) or ULTRASAT (Shvartzvald et al., 2024), the surrogates and accelerated inference techniques presented here could also be used to study different observing strategies to increase the efficiency of parameter estimation (Ragosta et al., 2024; Andrade et al., 2025) and maximize the science output of future detections.
@@ Acknowledgements.
We would like to thank Kaze Wong for useful discussions and assistance with the flowMC package. We would like to thank Nina Kunert for providing the lightcurve data of GRB211211A. We would like to thank Anna Neuweiler for her assistance with setting up the possis code. Computations were in part performed on the national supercomputer HPE Apollo Hawk at the High Performance Computing Center Stuttgart (HLRS) under the grant number GWanalysis/44189 and on the DFG-funded research cluster jarvis at the University of Potsdam (INST 336/173-1; project number: 502227537). This work used the Dutch national e-infrastructure with the support of the SURF Cooperative using grant no. EINF-8596. H. K., H. R., and T. D. acknowledge funding from the Daimler and Benz Foundation for the project “NUMANJI” and from the European Union (ERC, SMArt, 101076369). T.W. is supported by the research program of the Netherlands Organization for Scientific Research (NWO) under grant number OCENW.XL21.XL21.038. P.T.H.P. is supported by the research program of the Netherlands Organization for Scientific Research (NWO) under grant number VI.Veni.232.021. M.B. acknowledges the Department of Physics and Earth Science of the University of Ferrara for the financial support through the FIRD 2024 grant. Views and opinions expressed are those of the authors only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them.
@@ Appendix A GRB afterglow tophat jet surrogates
Figure 12: Benchmarks of the surrogate for the afterglowpy tophat jet model. We show the deviations of surrogate predictions against a test data set of size $n_{\text{test}}=7500$. Figure layout as in Fig. 1.
@@
Figure 13: Benchmarks of the surrogate for the pyblastafterglow tophat jet model. We show the deviations of surrogate predictions against a test data set of size $n_{\text{test}}=7680$. Figure layout as in Fig. 1.
@@
In this appendix, we present the benchmarks of the surrogates for the tophat jet models of afterglowpy and pyblastafterglow. Figure 12 and 13 show the distribution the prediction error of the surrogate against a test data set. The different architectures are shown again as different colors. In general, all trained surrogates perform well on the test data set, as the typical mean square error across the entire lightcurve is $\lesssim 0.1$ mag across all bandpasses. The only exception is the X-ray filter for pyblastafterglow, but this can be attributed again to the hard cut below $2\cdot 10^{-22}$ mJys we imposed on the training data (see Sec. 3.2).
@@
For afterglowpy, the MLP surrogate performs noticeably worse than the corresponding one with a cVAE architecture, especially at later times. Considering some example lightcurves, we find that the reason for this is likely due to the fact that the MLP architecture performs worse when a counterjet causes a late bump in the lightcurve, whereas the cVAE does not struggle with this feature as much. However, the counterjet contribution is also incorporated in the Gaussian models above and in the pyblastafterglow tophat model, where the MLP and cVAE perform similar. For both afterglowpy and pyblastafterglow, the absolute mismatch across the lightcurve of the cVAE surrogate is typically confined within 0.2 mag, except again in the X-ray filter for pyblastafterglow. For the afterglowpy cVAE, 95% of test predictions are always within 1 mag of the proper lightcurve, for the pyblastafterglow cVAE it is even 97%.
Appendix B Injection-recoveries for the pyblastafterglow and possis surrogates
@@
Figure 14: Parameter recovery of an injected mock lightcurve from the Gaussian pyblastafterglow jet model. Figure layout as in Fig. 4. The injection is from the pyblastafterglow base model, the recovery is with the fiesta surrogate and the flowMC sampler.
@@
Figure 15: Parameter recovery of an injected mock lightcurve from the Gaussian pyblastafterglow jet model. Figure layout as in Fig. 4. The injection is from the possis directly, the recovery is with the fiesta surrogate and the flowMC sampler. The insets on the upper right side show the injection data with the best-fit lightcurve (purple) and the true possis lightcurve (red). Triangles indicate upper mock detection limits at 24 mag placed to prevent the surrogate being used on parts of the lightcurve where the training data is subject to Monte-Carlo noise.
@@
In Fig. 14 we show the posterior corner plots from injections with the pyblastafterglow afterglow model. Likewise, in Fig. 15 we show the posterior from an injection with the possis KN model. The meaning of the respective parameters is given in Tab. 1, the priors are uniform in the ranges specified there. We set the prior for $\sigma_{\rm sys}$ uniform from 0.3 to 1 mag in case of the pyblastafterglow injection and uniform from 0.5 to 1 mag in case of the possis injection, as in the latter case the surrogate is expected to have slightly larger prediction errors. We find that the true injected values are well recovered in both cases. For the recovery of the pyblastafterglow injection we find that $\sigma_{\rm sys}$, sampled from a uniform prior $\mathcal{U}(0.3,1)$ converges to a value slightly larger than the lower prior bound. We attribute this to the late bump in the pyblastafterglow lightcurve at 100–150 days that is not accurately accounted for by the surrogate in the optical. On this interval, the surrogate deviates from the true lighcurve by about 0.4 mag, which the systematic uncertainty takes into account. However, this only works because the surrogate matches the other parts of the lightcurve within 0.3 mag and the larger deviation is confined to a small segment. In general, sampling $\sigma_{\rm sys}$ without a sufficiently high prior bound may fail to adequately capture the surrogate prediction error.

### v2

@@ 6 Discussion and conclusion
The fiesta package provides ML surrogates of KN and GRB afterglow models to enable efficient likelihood evaluation when sampling a posterior from photometric transient data. It offers a flexible API for training these surrogates, enabling the use of effectively three different architectures. Since fiesta utilizes the jax framework, training can be hardware-accelerated by execution on a GPU. Additionally, fiesta includes functionalities to sample the light curve posterior with flowMC (Wong:2022xvh), an adaptive MCMC sampler that uses gradient-based samplers, and normalizing flow proposals that can also be run on GPUs.
@@
To demonstrate fiesta’s features, we present surrogates for the afterglowpy (Ryan:2019fhz) and pyblastafterglow (Nedora:2024vrv) GRB afterglow models, as well as for the KN model from possis (Bulla:2019muo; Bulla:2022mwo). We find that the surrogates perform well against the respective test data sets, as the prediction error is usually bound within 0.3–0.5 mag. When used during inference, this finite prediction error needs to be offset with a systematic uncertainty, $\sigma_{\rm sys}$, which can also be sampled in a flexible manner within fiesta’s implementation. We applied these surrogates to two events with joint KN and GRB afterglow emission, namely AT2017gfo/GRB170817A and GRB211211A. Using our surrogates, we find similar results to previous analyses of these events, even when using the new pyblastafterglow model for the GRB afterglow. However, our posteriors can be evaluated within minutes; whereas previous analyses, for instance, those of Pang:2022rzc; Koehn:2024set, could take several hours to days, mainly due to the more expensive likelihood evaluation when computing the light curve directly from afterglowpy.
@@
Despite the advantages regarding speed when relying on ML surrogates in the likelihood evaluation, this approach also comes with some drawbacks. For one, the parameter ranges in which our surrogates interpolate the training data are fixed and, therefore, it is not possible to extend priors without retraining the surrogates. However, these surrogates can easily be fine-tuned on new training datasets for future use in other places in the parameter space. Further, the prediction error from the surrogate compared to the physical base model could introduce biases in the posteriors. We assessed these potential biases by creating afterglowpy injection recoveries, where the posterior can be obtained with the surrogate and with the base model. We find that the values for the KL divergence match theoretical expectations within the surrogate uncertainty (Bevins:2025srk). Furthermore, while we did find biases in the recovery of certain parameters (based on our P-P plots in Figs. 5 and 6), we do not find that this bias is necessarily caused by the use of our surrogates. Still, failure of the surrogate in certain parameter regions cannot be excluded and sanity checks are recommended to validate the results. For instance, the physical base model could be run with the best-fit parameters from the posterior to verify that they actually provide a good fit to the data.
@@
The use of ML techniques for accelerated likelihood evaluation and efficient sampling in Bayesian inference is a common practice in various areas of astrophysics. For instance, the analysis of X-ray spectra has recently been accelerated by the use of advanced stochastic samplers written in jax (jaxspec) or with the help of ML surrogate models (Barret:2024kvc; Tutone:2025hwh; Dupourque:2025vqe). For supernovae, Gaussian process surrogate models were developed for inference in Simongini:2024chh; Simongini:2025gse. Leeney:2025pwe created a jax-compatible version of sncosmo (sncosmo). In fiesta, we focused on surrogates for electromagnetic transients of BNS mergers; however, in principle, surrogates for other types of transients could also be incorporated. Additionally, as fiesta surrogates utilize the jax framework, this opens up the possibility for them to be combined with other jax software, such as jim (Wong:2023lgb; Wouters:2024oxj) for the parameter estimation on GW signals from BNSs. This would enable fast sampling for the joint analysis of GW signals and electromagnetic counterparts and could be linked further to the inference of the EOS for NS matter through the jester (Wouters:2025zju) package, which also leverages jax to efficiently evaluate the likelihood of an EOS candidate with a given set of NS observations. Additionally, the backward compatibility of our surrogates to nmma provides the opportunity to perform efficient Bayesian model comparison to either compare transient types (Breschi:2021tbm; Kunert:2023vqd; Hussenot-Desenonges:2023oll) or to investigate different systematics in the modeling. Such investigations could be related to KN ejecta morphology (Heinzel:2020qlt; King:2025tqo; Hussenot-Desenonges:2025gik), KN heating rates and opacities (Tanaka:2019iqp; Bulla:2022mwo; Sarin:2024tja; Brethauer:2024zxg), GRB jet structure (Kunert:2023vqd; Hayes:2019hso; Lin:2021auf) or environment (Aksulu:2021crt), or mappings from BNS parameters to ejecta parameters (Ristic:2025bvt). Given the launch of new transient observatories such as the Vera C. Rubin Observatory (Andreoni:2024pkp) or ULTRASAT (Shvartzvald:2023ofi), the surrogates and accelerated inference techniques presented here could also be used to study different observing strategies to increase the efficiency of parameter estimation (Ragosta:2024bfr; Andrade:2025gme) and maximize the science output of future detections.
@@ Acknowledgements.
We would like to thank the anonymous referee for their helpful suggestions. We would like to thank Kaze Wong for useful discussions and assistance with the flowMC package. We would like to thank Nina Kunert for providing the light curve data of GRB211211A. We would like to thank Anna Neuweiler for her assistance with setting up the possis code. Computations were in part performed on the national supercomputer HPE Apollo Hawk at the High Performance Computing Center Stuttgart (HLRS) under the grant number GWanalysis/44189 and on the DFG-funded research cluster jarvis at the University of Potsdam (INST 336/173-1; project number: 502227537). This work used the Dutch national e-infrastructure with the support of the SURF Cooperative using grant no. EINF-8596. H. K., H. R., and T. D. acknowledge funding from the Daimler and Benz Foundation for the project “NUMANJI” and from the European Union (ERC, SMArt, 101076369). T.W. is supported by the research program of the Netherlands Organization for Scientific Research (NWO) under grant number OCENW.XL21.XL21.038. P.T.H.P. is supported by the research program of the Netherlands Organization for Scientific Research (NWO) under grant number VI.Veni.232.021. M.B. acknowledges the Department of Physics and Earth Science of the University of Ferrara for the financial support through the FIRD 2024 grant. Views and opinions expressed are those of the authors only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them.
@@ Appendix A GRB afterglow tophat jet surrogates
Figure 12: Benchmarks of the two surrogates for the afterglowpy tophat jet model. We show the deviations of surrogate predictions against a test data set of size $n_{\text{test}}=7500$. Figure layout is the same as in Fig. 1.
@@
Figure 13: Benchmarks of the two surrogates for the pyblastafterglow tophat jet model. We show the deviations of surrogate predictions against a test data set of size $n_{\text{test}}=7680$. Figure layout is the same as in Fig. 1.
@@
In this appendix, we present the benchmarks of the surrogates for the tophat jet models of afterglowpy and pyblastafterglow. Figures 12 and 13 show the distribution of the prediction error of the surrogate against a test data set. The different architectures are shown again as different colors. In general, all trained surrogates perform well, as the typical mean squared error across the entire light curve is $\lesssim 0.1$ mag across all filters. The only exception is the X-ray filter for pyblastafterglow, but this can be attributed again to the hard cut below $2\cdot 10^{-22}$ mJy we imposed on the training data (see Sect. 3.2).
@@
For afterglowpy, the MLP surrogate performs noticeably worse than the surrogate with a cVAE architecture, especially at later times. Considering some example light curves, we find that the reason for this is likely due to the fact that the MLP architecture performs worse when a counterjet causes a late bump in the light curve, whereas the cVAE does not struggle with this feature as much. However, the counterjet contribution is also incorporated in the Gaussian models above and in the pyblastafterglow tophat model, where the MLP and cVAE perform similarly. For both afterglowpy and pyblastafterglow, the absolute mismatch across the light curve of the cVAE surrogate is typically confined within 0.2 mag, except again in the X-ray filter for pyblastafterglow. For the afterglowpy cVAE, 95% of test predictions are always within 1 mag of the proper light curve; for the pyblastafterglow cVAE it is even 97%.
Appendix B Injection recoveries for the pyblastafterglow and possis surrogates
@@
Figure 14: Parameter recovery for an injected mock light curve from the Gaussian pyblastafterglow jet model. Figure layout as in Fig. 4. The injection is from the pyblastafterglow base model, the recovery is with the fiesta surrogate and the flowMC sampler.
@@
Figure 15: Parameter recovery for an injected mock light curve from the Gaussian pyblastafterglow jet model. Figure layout as in Fig. 4. The injection is from the possis directly, the recovery is with the fiesta surrogate and the flowMC sampler. The insets on the upper right side show the injection data with the best-fit light curve (purple) and the true possis light curve (red). Triangles indicate upper mock detection limits at 24 mag placed to prevent the surrogate being used on parts of the light curve where the training data is subject to Monte Carlo noise.
@@
In Fig. 14, we show the posterior corner plots from an injection with the pyblastafterglow afterglow model. Likewise, in Fig. 15 we show the posterior from an injection with possis. The meaning of the respective parameters is given in Table 3, the priors are uniform within the ranges specified there. We set the prior for $\sigma_{\rm sys}$ to be uniform from 0.3 to 1 mag in case of the pyblastafterglow inference and uniform from 0.5 to 1 mag in case of the possis inference, since for the latter the surrogate is expected to have slightly larger prediction errors. We find that the true injected values are well recovered in both cases. For the recovery of the pyblastafterglow injection, we find that $\sigma_{\rm sys}$, sampled from a uniform prior $\mathcal{U}(0.3,1)$ converges to a value slightly larger than the lower prior bound. We attribute this to the late bump in the pyblastafterglow light curve at 100–150 days that is not accurately accounted for by the surrogate in the optical. On this interval, the surrogate deviates from the true light curve by about 0.4 mag, which the systematic uncertainty takes into account. However, this only works because the surrogate matches the other parts of the light curve within 0.3 mag and the larger deviation is confined to a small segment. In general, sampling $\sigma_{\rm sys}$ without a sufficiently high prior bound may fail to adequately capture the surrogate prediction error.